# Deep Evidential Learning for Dose Prediction

### v1

@@ Abstract
In this work, we present a novel application of an uncertainty-quantification framework called Deep Evidential Learning in the domain of radiotherapy dose prediction. Using medical images of the Open Knowledge-Based Planning Challenge dataset, we found that this model can be effectively harnessed to yield uncertainty estimates that inherited correlations with prediction errors upon completion of network training. This was achieved only after reformulating the original loss function for a stable implementation. We found that (i)epistemic uncertainty was highly correlated with prediction errors, with various association indices comparable or stronger than those for Monte-Carlo Dropout and Deep Ensemble methods, (ii)the median error varied with uncertainty threshold much more linearly for epistemic uncertainty in Deep Evidential Learning relative to these other two conventional frameworks, indicative of a more uniformly calibrated sensitivity to model errors, (iii)relative to epistemic uncertainty, aleatoric uncertainty demonstrated a more significant shift in its distribution in response to Gaussian noise added to CT intensity, compatible with its interpretation as reflecting data noise. Collectively, our results suggest that Deep Evidential Learning is a promising approach that can endow deep-learning models in radiotherapy dose prediction with statistical robustness. Towards enhancing its clinical relevance, we demonstrate how we can use such a model to construct the predicted Dose-Volume-Histograms’ confidence intervals.
@@ 1 Introduction
Radiotherapy treatment planning is a highly intricate process requiring collaborative interactions among medical physicists, dosimetrists and radiation oncologists in formulating the optimal treatment plan (see e.g. [1, 2]) for each individual patient. It is a procedure that is characterized by a potentially large degree of user-variability arising from different inter-institutional guidelines, the fundamental origin of which lies in the subjective nature of tradeoffs like those between tumor control and sparing of normal tissues, etc. [1, 3, 4] Automation by machine learning methods can help to homogenize these variations among the medical professionals, while attaining improvements in the consistency of overall plan quality (see e.g. [5]). The adoption of A.I.-related techniques led to what is known as knowledge-based planning (KBP) [6], which leverages knowledge inherent in past clinical treatment plans to generate new ones with minimal intervention from human experts.
@@
A complete KBP method can normally be regarded as a two-stage pipeline [7] : (i)prediction of the dose distribution that should be delivered to patient (ii)conversion of the prediction into a deliverable treatment plan via optimization. Recent dose prediction models cover quite a range of sites and modalities [5, 8], including prostate intensity-modulated-radiation-therapy (IMRT) [9, 10], prostate volumetric modulated arc therapy (VMAT) [11], lung IMRT [12] and head and neck VMAT [13]. These models were designed to predict volumetric dose distributions from which dose-volume histogram (DVH) and other dose constraint-related statistics can be deduced. Within the literature, it has been noted [6] that the vast majority of published works was performed using large private datasets which made comparison of model quality challenging. Towards alleviating this issue, the Open Knowledge-Based Planning (OpenKBP) Grand Challenge was organized [7, 14] to enable an international effort for the comparison of dose prediction models on a single open dataset involving 340 patients of head and neck cancer treated by IMRT. To date, there has been many excellent dose prediction models trained on the OpenKBP dataset as reviewed in [5]. Notable examples include DeepDoseNet of [15] which combined features of ResNet and DenseNet, the MtAA-NET of [16] which is based on a generative adversarial network, and TrDosePred of [17] and Swin UNETR++ of [18] which are Transformer-based frameworks.iiiFor a more complete list, see Table 5 of [5].. Like the above-mentioned projects, our work here leverages upon the OpenKBP dataset and attempts to furnish dose prediction. However, unlike the majority of these papers, the major focus of our work here lies in developing a dose prediction model that also encapsulates an uncertainty quantification framework.
@@
As we navigate towards integrating deep learning methods in the real clinic, a safety-related concern lies in whether and how the model can express its own uncertainty when making predictions based on real-life data distributions beyond those that they were trained on. Uncertainty quantification has been a major theme in the broader context of machine learning [19], and has recently gathered increasing attention in the domain of medical image analysis [20]. However, to our knowledge, there is a scarcity of works devoted specifically to uncertainty quantification in dose prediction models. As already noted in papers dedicated to medical image analysis [20, 21, 22, 23, 24], uncertainty estimates equip deep learning model with statistical robustness and a measure of reliability that can be useful in discovering regions of the model’s ignorance, thereby alerting the human expert to potentially erroneous predictions. In the domain of dose prediction, uncertainty estimates can be used to characterize the reliability of a model. In the KBP pipeline, dose prediction models are themselves inputs to some dose mimicking model [7] which generates deliverable treatments via optimization of a set of objective functions [25, 26]. Apart from using reliability as a selection criterion for various dose prediction models, it was pointed out in [7, 27] that probabilistic dose distributions can serve as robust inputs to the objective functions. At the time of writing, we are only aware of [8] and [28] being the only publications that examined the role of uncertainty quantification frameworks for dose prediction. The work in [8] studied how Monte-Carlo Dropout (MC Dropout) and a Deep Ensemble-based bagging method can furnish useful uncertainty estimates for a U-Net-based dose prediction model, whereas [28] proposed a Gaussian mixture model in which the standard deviation of each Gaussian mixture is part of the network’s outputs and can be used to reflect the tradeoff implicit in data from two different treatment protocols.
@@
In this work, we examined the applicability of a relatively new uncertainty quantification framework known as Deep Evidential Learning [29, 30] for dose prediction. Its fundamental principle lies in asserting a higher-order Bayesian prior over the probabilistic neural network output, embedding it within the loss function and subsequently generating uncertainty estimates together with the completion of model training. There are two major variants of it as proposed in the seminal works of Sensoy et al. in [30] [‘Deep Evidential Classification’] and Amini et al. in [29] [‘Deep Evidential Regression’]. We apply the formalism of [29] to construct a dose prediction model with 3D U-Net [31] as the backbone model architecture. In the process, we found that this approach required additional crucial refinements for an effective and stable implementation. These refinements are related to the form of the final model layer which connects inputs and weights to the parameters of the higher-order Bayesian prior distribution. These parameters are formulated as the model’s outputs, and are thus naturally obtained when training is completed. They are then used to furnish estimates of model uncertainty. This is in contrast to two major categories of uncertainty quantification frameworks in the current literatureiiiiiiA general survey can be found in [19] whereas comprehensive recent reviews in the field of medical imaging can be found in [20] and [32]. : (i) MC Dropout (ii) Deep Ensemble method. For MC Dropout [33], one typically inserts stochastic dropout variables at various points of the neural networks, and after training, a number of feedforward passes are performed to evaluate the mean and variance of the output. The latter is then regarded as the model uncertainty. For Deep Ensemble method [34, 35], one typically collects a family of models sharing the main network structure but differing in hyperparameters or simply the initial random weight distributions. From the ensemble set of outputs, the model prediction is taken as the average with uncertainty being the variance.
@@ 2.1 Model uncertainties from a Bayesian prior probability distribution
    Consider an ensemble of models, not necessarily identical in architecture, being trained upon the same dataset to yield estimates of the parameters $(\mu_{k},\sigma_{k})$, we would then obtain a spectrum of differing values of $(\mu_{k},\sigma_{k})$, the distribution being stochastic in nature due to the choice of model hyperparameters and their sensitivity to the random noise present in the training dataset.
@@
$\mu\sim\mathcal{N}(\gamma,\sigma^{2}/\nu),\,\,\,p(\mu|\gamma,\sigma^{2}/\nu)=\frac{\sqrt{\nu}}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{\nu(y-\mu)^{2}}{2\sigma^{2}}},$ (2)
@@ 2.2 Deep Evidential regression for a dose prediction model: a refined loss function
where the expression lying on the right of the term $1/(y(1-y))$ can be identified as the Student’s t distribution $\text{St}\left(\log(\frac{y}{1-y});\gamma,\frac{\beta(1+\nu)}{\nu\alpha},2\alpha\right)$. In [29], the model’s primary loss function was then derived from eqn. (8) by taking the negative logarithm of it and adding a suitable regularization loss term.
Unfortunately, we found that in this form, the loss function did not lead to stable learning curves that could generate uncertainty heatmaps that correlated with prediction errors. After some experimentation, we found that making the following refinements to eqn. (8) significantly enhanced the effectiveness of the model:
@@
$L_{pri}$ $=$ $f_{s}\left(\frac{\Gamma(\alpha+\frac{1}{2})}{\Gamma(\alpha)}\sqrt{\frac{\nu}{\pi}}\left[2\beta(1+\nu)\right]^{\alpha}\left[\nu\left(\log\left(\frac{y}{1-y}\right)-\gamma\right)^{2}+2\beta(1+\nu)\right]^{-(\alpha+\frac{1}{2})}\right),\,\,\,$ (9)
$\mathcal{L}_{f}$ $=$ $\mathcal{L}_{pri}+\lambda_{KL}|y-\gamma|(2\nu+\alpha)+\lambda_{mse}\left(\log\left(\frac{y}{1-y}\right)-\gamma\right)^{2},\,\,\,f_{s}(g)\equiv\frac{1}{1+g},$ (10)
where $\lambda_{KL},\lambda_{mse}$ are crucial hyperparameters of which optimal values for our purpose will be discussed in Sec. 3.2. The first regularization term $\mathcal{L}_{reg}=|y-\gamma|(2\nu+\alpha)$ preceded by $\lambda_{KL}$ was proposed in [29] to penalize statistical evidence for incorrectly labeled terms. We note in passing that in [29], it was argued that the prior (the function $\mathcal{P}(\mu,\sigma|\alpha,\beta,\nu,\gamma)$ in eqn. (4) ) has a singular limit when we take parameter values corresponding to ‘zero evidence’, and that softening this limit by introducing finite yet small values for $\alpha=1+\epsilon$, $\nu=\epsilon$ didn’t appear to be effective. We found this to be the case as well. Another possibility that we explored is using Jeffrey’s prior (which in this case involves an improper uniform prior for the mean of the model output). But this turned out to be similarly ineffective relative to the choice adopted in [29].
@@ 2.3 Related Work
The work in [28] proposes the use of Gaussian mixture models of which parameters are the neural network’s outputs, and with U-Net as the backbone architecture. Such a mixture density network is very similar in form to our Deep Evidential Learning model since the output variables of our model are also parameters of a PDF. In our case, the PDF is a higher-order normal-inverse-gamma distribution whereas in [28], the parameters are the means and variances of the component normal distributions (see eqn. (1)) and the coefficients defining their linear combination. Each component of the Gaussian mixture pertained to a treatment protocol representing specific priorities (e.g. organs at risk (OAR) sparing over target coverage, etc.), and the relative values of the variances quantified the trade-offs between these protocols. The cross-entropy loss function associated with the cumulative distribution function (CDF) of the predicted dose was then taken as part of the objective function for generating a dose mimicking model. In [28], deliverable treatment plans were the endpoints of their dose prediction algorithm, with the optimization problem solved by RayStation’s native sequential quadratic programming solver and final dose computed via a collapsed cone algorithm [28]. Although we noted that there was no scrutiny of the correlation between uncertainty and error, the authors of [28] (and its follow-up work [27]) notably pointed out that probabilistic dose prediction models such as [8] and ours could play a crucial role in their pipeline since our model would yield dose CDFs that can be used to define the cross-entropy loss function as part of the dose-mimicking model’s objective function. It would thus be interesting to extend results of our work here by integrating our probabilistic dose predictions into a pipeline like [28] where deliverable treatment plans with specified machine parameter settings can be anticipated as final outcomes.
@@ 3 Methodology
3.1 On the OpenKBP Challenge Datset
The OpenKBP Challenge dataset of [7, 14] is devoted towards establishing an open framework for the development of plan optimization models for knowledge-based planning (KBP) in radiotherapy. The data was obtained for 340 patients with head-and-neck cancer. The dataset split proposed for training deep learning models was: 200 (training), 40 (validation), 100 (testing). In our work, we followed this decomposition of data.
@@ 3.2 On model architecture and hyperparameters
Since our primary goal is to examine the uncertainty estimation aspects of the model, in this work, we adopt a simple vanilla 3D U-Net as the backbone architecture upon which we insert two additional layers so that they are compatible with the framework of Deep Evidential Learning.
@@
For the loss function originally proposed in [29], we found that unfortunately, it led to training curves with frequent oscillations and poor eventual outcomes. As explained in detail in Sec. 2.2, we reformulated the loss function for a logistic representation of the dose variable such that one could obtain a stable implementation. Our refined loss function in eqn. (10) is characterized by the hyperparameters $\lambda_{mse},\lambda_{KL}$ which describe couplings of the loss function to a mean-squared-error (MSE) term and a KL-divergence-like regularization term respectively. Results from ablation experiments with validation dataset indicated that optimal values of these hyperparameters lie within the intervals $\lambda_{mse}\in(0.01,0.1)$, $\lambda_{KL}\lesssim 0.01$. In particular, the finite interval of optimal $\lambda_{mse}$ arose from our finding that enhancing MSE-type performance too much would compromise the correlation between uncertainty and prediction error. For the results reported here for the rest of our paper, we took $\lambda_{mse}=0.05,\,\lambda_{KL}=0.01$. In Fig. 2, we plot the learning curves pertaining to the original and our refined loss functions for comparison.
@@
Towards comparing the Deep Evidential model against more common approaches of uncertainty quantification in literature, we implemented a Deep Ensemble model and Monte-Carlo Dropout model for the same dataset. Each of the 5 neural networks used in the ensemble model share the same 3D U-Net architecture with all hyperparameters preserved. Random weight initialization was all performed via the He-uniform initializer [39]. Similarly, for the Monte-Carlo Dropout model, the same base model was used for passing 30 forward passes with the dropout layers activated for model prediction. All models were trained for 200 epochs with a learning rate of $10^{-4}$ using Adam optimizer. Our focus in this work lies in studying aspects of uncertainty estimation, and thus our choice of a relatively simple backbone architecture equipped with only $\sim 6\times 10^{6}$ weight parameters. Compared to other more complicated setups, each of these models converged quickly in about a couple of hours while attaining a mean-absolute-error of dose prediction that laid within the top 20 of the score chart of the OpenKBP Challenge [14].

### v2

# Deep Evidential Learning for Radiotherapy Dose Prediction

@@ Abstract
Background: As we navigate towards integrating deep learning methods in the real clinic, a safety concern lies in whether and how the model can express its own uncertainty when making predictions. In this work, we present a novel application of an uncertainty-quantification framework called Deep Evidential Learning in the domain of radiotherapy dose prediction.
Method: Using medical images of the Open Knowledge-Based Planning Challenge dataset, we found that this model can be effectively harnessed to yield uncertainty estimates that inherited correlations with prediction errors upon completion of network training. This was achieved only after reformulating the original loss function for a stable implementation.
Results: We found that (i)epistemic uncertainty was highly correlated with prediction errors, with various association indices comparable or stronger than those for Monte-Carlo Dropout and Deep Ensemble methods, (ii)the median error varied with uncertainty threshold much more linearly for epistemic uncertainty in Deep Evidential Learning relative to these other two conventional frameworks, indicative of a more uniformly calibrated sensitivity to model errors, (iii)relative to epistemic uncertainty, aleatoric uncertainty demonstrated a more significant shift in its distribution in response to Gaussian noise added to CT intensity, compatible with its interpretation as reflecting data noise.
Conclusion: Collectively, our results suggest that Deep Evidential Learning is a promising approach that can endow deep-learning models in radiotherapy dose prediction with statistical robustness. We have also demonstrated how this framework leads to uncertainty heatmaps that correlate strongly with model errors, and how it can be used to equip the predicted Dose-Volume-Histograms with confidence intervals.
keywords:
uncertainty quantification , deep evidential learning , radiotherapy dose prediction
[…]
[inst1]Gryphon Center for Artificial Intelligence and Theoretical Sciences, country=Singapore
[…]
[inst2]Georgia Institute of Technology, Atlanta, GA, country=USA
[…]
[inst3]University of Pennsylvania, Perelman School of Medicine, Department of Radiation Oncology, Philadelphia, country=USA
@@ 1 Introduction
Radiotherapy treatment planning is a highly intricate process requiring collaborative interactions among medical physicists, dosimetrists and radiation oncologists in formulating the optimal treatment plan (see e.g. [1, 2]) for each individual patient. It is a procedure that is characterized by a potentially large degree of user-variability arising from different inter-institutional guidelines, the fundamental origin of which lies in the subjective nature of tradeoffs like those between tumor control and sparing of normal tissues, etc. [1, 3, 4] Automation by machine learning methods can help to homogenize these variations among the medical professionals, while attaining improvements in the consistency of overall plan quality (see e.g. [5]). The adoption of artificial intelligence (A.I.) related techniques led to what is known as knowledge-based planning (KBP) [6], which leverages knowledge inherent in past clinical treatment plans to generate new ones with minimal intervention from human experts.
@@
A complete KBP method can normally be regarded as a two-stage pipeline [7] : (i)prediction of the dose distribution that should be delivered to patient (ii)conversion of the prediction into a deliverable treatment plan via optimization. Recent dose prediction models cover quite a range of sites and modalities [5, 8], including prostate intensity-modulated-radiation-therapy (IMRT) [9, 10], prostate volumetric modulated arc therapy (VMAT) [11], lung IMRT [12] and head and neck VMAT [13]. These models are designed to predict volumetric dose distributions from which dose-volume histogram (DVH) and other dose constraint-related statistics can be deduced. Within the literature, it has been noted [6] that the vast majority of published works was performed using large private datasets which made comparison of model quality challenging. Towards alleviating this issue, the Open Knowledge-Based Planning (OpenKBP) Grand Challenge was organized [7, 14] to enable an international effort for the comparison of dose prediction models on a single open dataset involving 340 patients of head and neck cancer treated by IMRT. To date, there has been many excellent dose prediction models trained on the OpenKBP dataset as reviewed in [5]. Notable examples include DeepDoseNet of [15] which combined features of ResNet and DenseNet, the MtAA-NET of [16] which is based on a generative adversarial network, and TrDosePred of [17] and Swin UNETR++ of [18] which are Transformer-based frameworks.iiiiiiFor a more complete list, see Table 5 of [5].. Like the above-mentioned projects, our work here leverages upon the OpenKBP dataset and attempts to furnish dose prediction. However, unlike the majority of these papers, the major focus of our work here lies in developing a dose prediction model that also encapsulates an uncertainty quantification framework.
@@
As we navigate towards integrating deep learning methods in the real clinic, a safety-related concern lies in whether and how the model can express its own uncertainty when making predictions based on real-life data distributions beyond those that they were trained on. Uncertainty quantification has been a major theme in the broader context of machine learning [19], and has recently gathered increasing attention in the domain of medical image analysis [20]. However, to our knowledge, there is a scarcity of works devoted specifically to uncertainty quantification in dose prediction models. As already noted in papers dedicated to medical image analysis [20, 21, 22, 23, 24], uncertainty estimates equip deep learning models with statistical robustness and a measure of reliability that can be useful in discovering regions of the model’s ignorance, thereby alerting the human expert to potentially erroneous predictions. In the domain of dose prediction, uncertainty estimates can be used to characterize the reliability of a model. In the KBP pipeline, dose prediction models are themselves inputs to some dose mimicking model [7] which generates deliverable treatments via optimization of a set of objective functions [25, 26]. Apart from using reliability as a selection criterion for various dose prediction models, it was pointed out in [7, 27] that probabilistic dose distributions can serve as robust inputs to the objective functions. At the time of writing, we are only aware of [8] and [28] being the only publications that examined the role of uncertainty quantification frameworks for dose prediction. The work in [8] studied how Monte-Carlo Dropout (MC Dropout) and a Deep Ensemble-based bagging method can furnish useful uncertainty estimates for a U-Net-based dose prediction model, whereas [28] proposed a Gaussian mixture model in which the standard deviation of each Gaussian mixture is part of the network’s outputs and can be used to reflect the tradeoff implicit in data from two different treatment protocols.
@@
In this work, we examined the applicability of a relatively new uncertainty quantification framework known as Deep Evidential Learning [29, 30] for dose prediction. Its fundamental principle lies in asserting a higher-order Bayesian prior over the probabilistic neural network output, embedding it within the loss function and subsequently generating uncertainty estimates together with the completion of model training. There are two major variants of it as proposed in the seminal works of Sensoy et al. in [30] [‘Deep Evidential Classification’] and Amini et al. in [29] [‘Deep Evidential Regression’]. We apply the formalism of [29] to construct a dose prediction model with 3D U-Net [31] as the backbone model architecture. In the process, we found that this approach required additional crucial refinements for an effective and stable implementation. These refinements are related to the form of the final model layer which connects inputs and weights to the parameters of the higher-order Bayesian prior distribution. These parameters are formulated as the model’s outputs, and are thus naturally obtained when training is completed. They are then used to furnish estimates of model uncertainty. This is in contrast to two major categories of uncertainty quantification frameworks in the current literatureiiiiiiiiiA general survey can be found in [19] whereas comprehensive recent reviews in the field of medical imaging can be found in [20] and [32]. : (i) MC Dropout (ii) Deep Ensemble method. For MC Dropout [33], one typically inserts stochastic dropout variables at various points of the neural networks, and after training, a number of feedforward passes are performed to evaluate the mean and variance of the output. The latter is then regarded as the model uncertainty. For Deep Ensemble method [34, 35], one typically collects a family of models sharing the main network structure but differing in hyperparameters or simply the initial random weight distributions. From the ensemble set of outputs, the model prediction is taken as the average with uncertainty being the variance.
@@ 2.1 Model uncertainties from a Bayesian prior probability distribution
    For an ensemble of models being trained upon the same dataset to yield estimates of the parameters $(\mu_{k},\sigma_{k})$, one would obtain a spectrum of differing values of $(\mu_{k},\sigma_{k})$, the distribution being stochastic in nature due to distinct initial weight distributions, and their sensitivity to the random noise present in the training dataset.
@@
$\mu\sim\mathcal{N}(\gamma,\sigma^{2}/\nu),\,\,\,p(\mu|\gamma,\sigma^{2}/\nu)=\frac{\sqrt{\nu}}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{\nu(\gamma-\mu)^{2}}{2\sigma^{2}}},$ (2)
@@ 2.2 Deep Evidential Regression for a dose prediction model: a refined loss function
where the expression lying on the right of the term $1/(y(1-y))$ can be identified as the Student’s t distribution $\text{St}\left(\log(\frac{y}{1-y});\gamma,\frac{\beta(1+\nu)}{\nu\alpha},2\alpha\right)$. Unfortunately, we found that in this form, the loss function did not lead to stable learning curves that could generate uncertainty heatmaps that correlated with prediction errors. After some experimentation, we found that making the following refinements to eqn. (8) significantly enhanced the effectiveness of the model:
@@
$\mathcal{L}_{pri}$ $=$ $f_{s}\left(\frac{\Gamma(\alpha+\frac{1}{2})}{\Gamma(\alpha)}\sqrt{\frac{\nu}{\pi}}\left[2\beta(1+\nu)\right]^{\alpha}\left[\nu\left(\log\left(\frac{y}{1-y}\right)-\gamma\right)^{2}+2\beta(1+\nu)\right]^{-(\alpha+\frac{1}{2})}\right),\,\,\,$ (9)
$\mathcal{L}_{f}$ $=$ $\mathcal{L}_{pri}+\lambda_{KL}\left|\log\left(\frac{y}{1-y}\right)-\gamma\right|(2\nu+\alpha)+\lambda_{mse}\left(\log\left(\frac{y}{1-y}\right)-\gamma\right)^{2},\,f_{s}(g)\equiv\frac{1}{1+g}$ (10)
where $\lambda_{KL},\lambda_{mse}$ are crucial hyperparameters of which optimal values for our purpose will be discussed in Sec. 3.2. The first regularization term $\mathcal{L}_{reg}=\left|\log\left(\frac{y}{1-y}\right)-\gamma\right|(2\nu+\alpha)$ preceded by $\lambda_{KL}$ was proposed in [29] to penalize statistical evidence for incorrectly labeled terms. We note in passing that in [29], it was argued that the prior (the function $\mathcal{P}(\mu,\sigma|\alpha,\beta,\nu,\gamma)$ in eqn. (4) ) has a singular limit when we take parameter values corresponding to ‘zero evidence’, and that softening this limit by introducing finite yet small values for $\alpha=1+\epsilon$, $\nu=\epsilon$ did not appear to be effective. We found this to be the case as well. Another possibility that we explored is using Jeffrey’s prior (which in this case involves an improper uniform prior for the mean of the model output). But this turned out to be similarly ineffective relative to the choice adopted in [29].
@@ 2.3 Related Work
The work in [28] proposes the use of Gaussian mixture models of which parameters are the neural network’s outputs, and with U-Net as the backbone architecture. Such a mixture density network is very similar in form to our Deep Evidential Learning model since the output variables of our model are also parameters of a PDF. In our case, the PDF is a higher-order normal-inverse-gamma distribution whereas in [28], the parameters are the means and variances of the component normal distributions (see eqn. (1)) and the coefficients defining their linear combination. Each component of the Gaussian mixture pertained to a treatment protocol representing specific priorities (e.g. OAR sparing over target coverage, etc.), and the relative values of the variances quantified the trade-offs between these protocols. The cross-entropy loss function associated with the cumulative distribution function (CDF) of the predicted dose was then taken as part of the objective function for generating a dose mimicking model. In [28], deliverable treatment plans were the endpoints of their dose prediction algorithm, with the optimization problem solved by RayStation’s native sequential quadratic programming solver and final dose computed via a collapsed cone algorithm [28]. Although we noted that there was no scrutiny of the correlation between uncertainty and error, the authors of [28] (and its follow-up work [27]) notably pointed out that probabilistic dose prediction models such as [8] and ours could play a crucial role in their pipeline since our model would yield dose CDFs that can be used to define the cross-entropy loss function as part of the dose-mimicking model’s objective function. It would thus be interesting to extend results of our work here by integrating our probabilistic dose predictions into a pipeline like [28] where deliverable treatment plans with specified machine parameter settings can be anticipated as final outcomes.
@@ 3 Methodology
3.1 On the OpenKBP Challenge Dataset
The OpenKBP Challenge dataset of [7, 14] is devoted towards establishing an open framework for the development of plan optimization models for knowledge-based planning (KBP) in radiotherapy. It is an augmented variation of real clinical data constructed as follows: medical images were taken from several institutions available on The Cancer Imaging Archive (TCIA) [37] which hosts open-source data that has undergone de-identification compliant with DICOM standard.ivivivIn particular, all metadata and Protected Health Information (PHI) has been removed from the datasets provided in OpenKBP Challenge [7, 14]. After removing highly incomplete imaging datasets to obtain a final competition dataset comprising of 340 patients, synthetic radiation plans were generated for each of them using a published automated planning method [38]. This final dataset is then split randomly by the organizers of OpenKBP Challenge to be as follows: 200 (training), 40 (validation), 100 (testing). In our work, we followed this decomposition of data. As in a conventional machine learning pipeline, the validation dataset was used for tuning of hyperparameters whereas the testing dataset was used for reporting various results in our work. We refer the reader to [39, 40] for more pedagogical expositions of data-splitting practices and algorithms.
@@ 3.2 On model architecture and hyperparameters
Since our primary goal is to examine the uncertainty estimation aspects of the model, in this work, we adopt a simple vanilla 3D U-Net as the backbone architecture upon which we insert two additional layers so that they are compatible with the framework of Deep Evidential Learning. U-Nets and their variants have featured heavily in many deep-learning-guided tasks related to radiotherapy dose prediction [8, 14]. They are essentially convolutional neural networks equipped with an encoder-decoder architecture that allows effective learning of both image features at various levels of resolution (see e.g. [43, 44] for reviews on U-Net and [45, 46] for pedagogical introductions to convolutional neural networks).
@@
For the loss function originally proposed in [29], we found that unfortunately, it led to training curves with frequent oscillations and poor eventual outcomes. As explained in detail in Sec. 2.2, we reformulated the loss function for a logistic representation of the dose variable such that one could obtain a stable implementation. Our refined loss function in eqn. (10) is characterized by the hyperparameters $\lambda_{mse},\lambda_{KL}$ which describe couplings of the loss function to a mean-squared-error (MSE) term and a KL-divergence-like regularization term respectively. Results from ablation experiments with validation dataset indicated that optimal values of these hyperparameters lie within the intervals $\lambda_{mse}\in(0.01,0.1)$, $\lambda_{KL}\lesssim 0.01$. For the results reported here for the rest of our paper, we took $\lambda_{mse}=0.05,\,\lambda_{KL}=0.01$. In Fig. 2, we plot the learning curves pertaining to the original and our refined loss functions for comparison.
@@
Towards comparing the Deep Evidential model against more common approaches of uncertainty quantification in literature, we implemented a Deep Ensemble model and Monte-Carlo Dropout model for the same dataset. Each of the 5 neural networks used in the ensemble model shared the same 3D U-Net architecture with all hyperparameters preserved. Random weight initialization was all performed via the He-uniform initializer [47]. Similarly, for the Monte-Carlo Dropout model, the same base model was used for passing 30 forward passes with the dropout layers activated for model prediction. All models were trained for 200 epochs with a learning rate of $10^{-4}$ using Adam optimizer. Our focus in this work lies in studying aspects of uncertainty estimation, and thus our choice of a relatively simple backbone architecture equipped with only $\sim 6\times 10^{6}$ weight parameters. Compared to other more complicated setups, each of these models converged quickly in about a couple of hours while attaining a mean-absolute-error of dose prediction that laid within the top 20 of the score chart of the OpenKBP Challenge [14].